How do I write a quadratic equation using -6, and -4 as solutions.
Answer by algebrahouse.com(1659)

"How do I write a quadratic equation using -6, and -4 as solutions."

Start with solution and work backwards:

x = -6 or x = -4
x + 6 = 0 or x + 4 = 0 {added 6 and added 4 to both sides}
(x + 6)(x + 4) = 0 {multiplied the two binomials and set equal to 0}
x² + 10x + 24 = 0 {used foil method}
www.algebrahouse.com